# Spectral representations of interpolation spaces of reproducing kernel Hilbert spaces

### v1

@@ 5 Proofs
As the RKHS $H$ has a bounded kernel, the eigenfunctions $e_{i}$ are bounded as well. Hence by Theorem 5.7, the lifted eigenfunctions $\overline{e}_{i}=\phi([e_{i}]_{\sim})$ are measurable, bounded and we know that for each $i\in I$, there is a $\nu$-null set $N_{i}$ such that $e_{i}(x)=\overline{e}_{i}(x)$ for all $x\in X\setminus N_{i}$. As $I$ is at most countable, $N:=\bigcup_{i\in I}N_{i}$ is again measurable with $\nu(N)=0$. By that, for all $x\in X\setminus N$ and all $i\in I$, we have $\overline{e}_{i}(x)=e_{i}(x)$. Furthermore, let us denote
$\kappa:=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (81)
(i). Let $b\in\ell^{(\mu,\theta,r)}$. Then by Lemma 5.6, we know that $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in $H_{\nu}^{(\theta,r)}$. Due to the continuity of the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$, it converges in $\mathrm{L}^{\infty}(\nu)$ as well and we have
$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (82)
By continuity of the lifting operator $\phi:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$ from Theorem 5.7, we thus find
$\phi\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)=\sum_{i\in I}b_{i}\phi([e_{i}]_{\sim})=\sum_{i\in I}b_{i}\overline{e}_{i}\text{,}$ (83)
@@
(iii). As $\ell^{(\mu,\theta,r)}$ is a Banach space by Remark 5.8, it follows that $\overline{H}_{\nu}^{(\theta,r)}$ is a Banach space as well. By the uniform convergence in (i), $\overline{H}_{\nu}^{(\theta,r)}$ consists of functions. It remains to show that the evaluation functionals are continuous. To this end, let us fix a $b\in\ell^{(\mu,\theta,r)}$. Because $\phi$ has operator norm at most $1$, by (83) we find that
$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}=\left\|\phi\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)\right\|_{\mathcal{L}^{\infty}(X)}\leq\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\text{.}$ (84)
Now for all $x\in X$, we consider $\delta_{x}:\mathcal{L}^{\infty}(X)\rightarrow\mathbb{R},f\mapsto f(x)$ to find that
$\left|\delta_{x}\left(\sum_{i\in I}b_{i}\overline{e}_{i}\right)\right|=\left|\sum_{i\in I}b_{i}\overline{e}_{i}(x)\right|\leq\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}$ (85)
where for second to last inequality, we combined (82) and (84).
(iv). This follows as the norms on $H_{\nu}^{(\theta,r)}$ and on $\overline{H}_{\nu}^{(\theta,r)}$ are both inherited from the $\ell^{(\mu,\theta,r)}$-norm of the unique Fourier coefficients.
(v). Note that by using (85), we find that
$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}=\kappa\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{H_{\nu}^{(\theta,r)}}\text{,}$ (86)
which shows $\leq$. For $\geq$, we note that
$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|f\|_{\mathcal{L}^{\infty}(X)}\geq\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|$ (87)
where the last step follows as $\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}=\|[f]_{\sim}\|_{H_{\nu}^{(\theta,r)}}$. ∎
@@
The assertion follows directly from Lemma 5.11 by using the inverse of the isometric isomorphism from (iv). ∎
@@
First of all, note that by Definition 2.2, an arbitrary $g]_{\sim}\in H_{\nu}^{(\theta,r)}$ can be written as a series
@@
for a suitable $b\in\ell^{(\mu,\theta,r)}$, where the convergence is in $H_{\nu}^{(\theta,r)}$ by Lemma 5.6. In addition, it holds that
$\|[g]_{\sim}\|_{H_{\nu}^{(\theta,r)}}=\|b\|_{(\mu,\theta,r)}\text{.}$ (89)
@@
“(i) $\Rightarrow$ (ii)”: By Property (i), there is a $\nu$-null set $N$ such that
@@
$\sum_{i\in I}|b_{i}e_{i}(x)|$ $=\sum_{i\in I}|b_{i}f_{i}(x)|\mu_{i}^{-\theta}$ (91)
$\leq\|(f_{i}(x))_{i\in I}\|_{(\mu,\theta,r^{\prime})}\|b\|_{(\mu,\theta,r)}$ (92)
$\leq\kappa\,\|b\|_{(\mu,\theta,r)}\text{,}$ (93)
@@
Consequently, we find for such an arbitrary $g]_{\sim}\in H_{\nu}^{(\theta,r)}$ that
@@
$=\kappa\|[g]_{\sim}\|_{H_{\nu}^{(\theta,r)}}$ (98)
by an application of (93). This proves (ii).
“(ii) $\Rightarrow$ (i)”: By Lemma 5.11, for almost all $x\in X$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
$\sum_{i\in I}b_{i}f_{i}(x)\mu_{i}^{-\theta}=\sum_{i\in I}b_{i}e_{i}(x)=\sum_{i\in I}b_{i}\overline{e}_{i}(x)\leq\kappa\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\overline{H}_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (99)
Thus, for almost all $x\in X$, an application of Lemma 5.10 to $a_{i}=f_{i}(x)$ yields that $(f_{i}(x))_{i\in I}\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|f_{i}(x)\|_{(\mu,\theta,r^{\prime})}\leq\kappa$. This finally shows (i). ∎
@@
Then by the norm equivalence from Theorem 3.1, we find a $C_{6}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.12:
$\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{6}\|(1-\mu_{i}g_{\lambda}(\mu_{i}))b_{i}\|_{(\mu,\theta,r)}$ (110)
$\leq C_{6}\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (111)
@@
(ii). By the norm equivalence in Theorem 3.1, we find a $C_{5}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.12:
$\|[f_{P,\lambda,g}]_{\sim}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{5}\|\mu_{i}g_{\lambda}(\mu_{i})b_{i}\|_{(\mu,\theta,r)}$ (112)
$\leq C_{5}\sup_{i\in I}\left(|\mu_{i}g_{\lambda}(\mu_{i})|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (113)

### v2

@@ 5 Proofs
[…]
Let us consider the $\nu$-null set $N:=N^{\prime}\cup\bigcup_{i\in I}N^{\prime\prime}_{i}$. Then for all $i\in I$, it holds that $\overline{e}_{i}:=\mathbbm{1}_{X\setminus N}\breve{e}_{i}$ is $\mathcal{A}$-measurable with $\overline{e}_{i}(x)=e_{i}(x)$ for all $x\in X\setminus N$.
Let us denote the linear operator
[…]
which is well-defined by Lemma 5.11 and has operator norm at most $1$ due to Lemma 5.11 and Theorem 5.7. Note that $A[e_{i}]_{\sim}=\overline{e}_{i}$ are $\mathcal{A}$-measurable functions on $X$.
Furthermore, we write
$\kappa:=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (99)
(i). Let $b\in\ell^{(\mu,\theta,r)}$. Then by Lemma 5.6, we know that $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in $H]_{\sim}^{(\theta,r)}$. Due to the continuity of the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$, it also converges in $\mathrm{L}^{\infty}(\nu)$, and we have
$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (100)
By continuity of the operator $A:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$, we thus find
$A\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)=\sum_{i\in I}b_{i}A([e_{i}]_{\sim})=\sum_{i\in I}b_{i}\overline{e}_{i}\text{,}$ (101)
@@
(iii). As $\ell^{(\mu,\theta,r)}$ is a Banach space by Remark 5.8, it follows that $\overline{H}_{\nu}^{(\theta,r)}$ is a Banach space as well. By the uniform convergence in (i), $\overline{H}_{\nu}^{(\theta,r)}$ consists of functions. It remains to show that the evaluation functionals are continuous. To this end, let us fix a $b\in\ell^{(\mu,\theta,r)}$. Because $A$ has operator norm at most $1$, we find by (101) that
$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}=\left\|A\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)\right\|_{\mathcal{L}^{\infty}(X)}\leq\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\text{.}$ (102)
Now let $\delta_{x}:\mathcal{L}^{\infty}(X)\rightarrow\mathbb{R},f\mapsto f(x)$ be the evaluation functional at some fixed $x\in X$. Combining (100) and (102), we find
$\left|\delta_{x}\left(\sum_{i\in I}b_{i}\overline{e}_{i}\right)\right|=\left|\left(\sum_{i\in I}b_{i}\overline{e}_{i}\right)(x)\right|\leq\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (103)
(iv). This follows as the norms on $H]_{\sim}^{(\theta,r)}$ and on $\overline{H}_{\nu}^{(\theta,r)}$ are both inherited from the $\ell^{(\mu,\theta,r)}$-norm of the unique Fourier coefficients.
(v). Note that by using (100) and (102), we find
$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}=\kappa\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{[H]_{\sim}^{(\theta,r)}}\text{,}$ (104)
which shows $\leq$. For $\geq$, we use
$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|f\|_{\mathcal{L}^{\infty}(X)}\geq\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|$ (105)
where the last step follows by $\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}=\|[f]_{\sim}\|_{[H]_{\sim}^{(\theta,r)}}$. ∎
@@
The assertion follows directly from Lemma 5.12 by using the inverse of the isometric isomorphism from (iv). The additional statement follows from (ii). ∎
In Theorem 3.2, we describe by means of lifting theory how to find representatives of $H]_{\sim}^{(\theta,r)}$ that give rise to a Banach space of functions. However, note that we already chose a set of “canonical” representatives in Definition 2.3. Let us analyse under which circumstances this set of “canonical” representatives forms a BSF. To this end, let us apply a topological technique similar to [23].
@@
Let us denote
$\kappa:=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (106)
(i). Let $b\in\ell^{(\mu,\theta,r)}$. Then by Lemma 5.6, we know that $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in $H]_{\sim}^{(\theta,r)}$. Due to the continuity of the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$, it converges in $\mathrm{L}^{\infty}(\nu)$ as well, and we have
$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (107)
Now the partial sums of $\sum_{i\in I}b_{i}e_{i}$ are $\tau(H)$-continuous and $\sum_{i\in I}b_{i}[e_{i}]_{\sim}=[\sum_{i\in I}b_{i}e_{i}]_{\sim}$ converges in $\mathrm{L}^{\infty}(\nu)$-norm. By Lemma C.4, we hence find that $\sum_{i\in I}b_{i}e_{i}$ converges in $C_{b}(X,\tau(H))$, hence uniformly.
(ii). First of all, note that the coefficients $b_{i}$ are uniquely determined by the function $\sum_{i\in I}b_{i}e_{i}$, as the $e_{i}]_{\sim}$ are $\mathrm{L}^{2}(\nu)$-orthogonal.
As $\ell^{(\mu,\theta,r)}$ is a Banach space by Remark 5.8, it follows that ${H}_{\nu}^{(\theta,r)}$ is a Banach space as well. By the uniform convergence in (i), ${H}_{\nu}^{(\theta,r)}$ consists of functions. It remains to show that the evaluation functionals are continuous. To this end, let us fix a $b\in\ell^{(\mu,\theta,r)}$. Then for all $x\in X$, we consider $\delta_{x}:\mathcal{L}^{\infty}(X)\rightarrow\mathbb{R},f\mapsto f(x)$ to find that
$\begin{split}\left|\delta_{x}\left(\sum_{i\in I}b_{i}e_{i}\right)\right|=\left|\left(\sum_{i\in I}b_{i}e_{i}\right)(x)\right|\leq\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{\mathcal{L}^{\infty}(X)}=\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\|b\|_{(\mu,\theta,r)}\text{,}\end{split}$ (108)
where for second to last step, we used Corollary C.3, and the last inequality is due to (107).
(iii). This follows as the norms on $H]_{\sim}^{(\theta,r)}$ and on ${H}_{\nu}^{(\theta,r)}$ are both inherited from the $\ell^{(\mu,\theta,r)}$-norm of the unique Fourier coefficients.
(iv). In the proof of (i), we found that ${H}_{\nu}^{(\theta,r)}$ consists of $\tau(H)$-continuous and bounded functions.
Note that by using the isometric isomorphism $\alpha:{H}_{\nu}^{(\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)}$ from (iii) and the isometry $\beta:\mathrm{C}_{b}(X,\tau(H))\rightarrow\mathrm{L}^{\infty}(\nu),f\mapsto[f]_{\sim}$, the following diagram commutes:
${H}_{\nu}^{(\theta,r)}$$\mathrm{C}_{b}(X,\tau(H))$$H]_{\sim}^{(\theta,r)}$$\mathrm{L}^{\infty}(\nu)$$\scriptstyle{\mathrm{Id}}\scriptstyle{\simeq}$$\scriptstyle{\alpha}$$\scriptstyle{\beta}$$\scriptstyle{\mathrm{Id}}$
But as $\mathrm{Id}:[H]_{\sim}^{(\theta,r)}\rightarrow\mathrm{L}^{\infty}(\nu)$ is bounded with operator norm $\kappa$, it follows that $\mathrm{Id}:{H}_{\nu}^{(\theta,r)}\rightarrow\mathrm{C}_{b}(X,\tau(H))$ is bounded with operator norm $\kappa$ as well.
(v). Let $f\in H$. Then by Lemma 5.3, there is a sequence $(b_{i})_{i\in I}\in\ell^{2}(\mu^{-1})$ such that $f]_{\sim}=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$. It holds that
$\ell^{2}(\mu^{-1})\hookrightarrow[\ell^{2}(I),\ell^{2}(\mu^{-1})]_{(\theta,r)}\triangleq\ell^{(\mu,\theta,r)}\text{,}$ (109)
where for the first step, we apply Lemma A.2 and for the second, we apply Lemma 5.2 to the counting measure on $I$ and the weight sequence $\mu^{-1}$ to find that the spaces $\ell^{2}(I),\ell^{2}(\mu^{-1})]_{(\theta,r)}$ and $\ell^{(\mu,\theta,r)}$ are norm equivalent.
We hence find that $(b_{i})_{i\in I}\in\ell^{(\mu,\theta,r)}$. Now by (i), the sum $\sum_{i\in I}b_{i}e_{i}$ converges uniformly in $C_{b}(X,\tau(H))$. Because $f$ itself is $\tau(H)$-continuous as well, we have $f=\sum_{i\in I}b_{i}e_{i}$, as almost everywhere agreeing $\tau(H)$-continuous functions agree everywhere by Corollary C.3. This shows that $f=\sum_{i\in I}b_{i}e_{i}\in H^{(\theta,r)}$. ∎
Proof of Theorem 3.4.
First of all, note that by Definition 2.3, an arbitrary $g]_{\sim}\in[H]_{\sim}^{(\theta,r)}$ can be written as a series
@@
for a suitable $b\in\ell^{(\mu,\theta,r)}$, where the convergence is in $H]_{\sim}^{(\theta,r)}$ by Lemma 5.6. In addition, it holds that
$\|[g]_{\sim}\|_{[H]_{\sim}^{(\theta,r)}}=\|b\|_{(\mu,\theta,r)}\text{.}$ (111)
@@
Concerning the different assumptions, note that Assumptions H+ and T imply Assumption H.
“(i) $\Rightarrow$ (ii)” Obvious.
“(ii) $\Rightarrow$ (iii)” By Property (ii), there is a $\nu$-null set $N$ such that
@@
$\begin{split}\sum_{i\in I}|b_{i}e_{i}(x)|&=\sum_{i\in I}|b_{i}f_{i}(x)|\mu_{i}^{-\theta}\\ &\leq\|(f_{i}(x))_{i\in I}\|_{(\mu,\theta,r^{\prime})}\|b\|_{(\mu,\theta,r)}\\ &\leq\kappa\,\|b\|_{(\mu,\theta,r)}\text{,}\end{split}$ (113)
@@
$=\kappa\|[g]_{\sim}\|_{[H]_{\sim}^{(\theta,r)}}$ (118)
by an application of (113). This proves (iii).
“(iii) $\Rightarrow$ (ii)” under Assumption H+: By Lemma 5.12, for almost all $x\in X$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
$\left|\sum_{i\in I}b_{i}f_{i}(x)\mu_{i}^{-\theta}\right|=\left|\sum_{i\in I}b_{i}e_{i}(x)\right|=\left|\sum_{i\in I}b_{i}\overline{e}_{i}(x)\right|\leq\kappa\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\overline{H}_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{, }$ (119)
where $\overline{e}_{i}$ is defined as in Lemma 5.12. Note that the sums are absolutely convergent by Remark B.2. For almost all $x\in X$, an application of Lemma 5.10 to $a_{i}:=f_{i}(x)$ yields that $(f_{i}(x))_{i\in I}\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|f_{i}(x)\|_{(\mu,\theta,r^{\prime})}\leq\kappa$. This shows (ii).
“(ii) $\Rightarrow$ (i)” under Assumption T: An application of Lemma C.1 shows that the function $X\rightarrow[0,\infty]$ given by
$x\mapsto\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}$ (120)
is non-negative and lower semicontinuous. Hence by Lemma C.2, the essential supremum and the supremum agree.
“(iii) $\Rightarrow$ (ii)” under Assumption T: By Theorem 3.3, for all $x\in X$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
$\left|\sum_{i\in I}b_{i}f_{i}(x)\mu_{i}^{-\theta}\right|=\left|\sum_{i\in I}b_{i}e_{i}(x)\right|\leq\kappa\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{H_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (121)
Thus, for all $x\in X$, another application of Lemma 5.10 to $a_{i}:=f_{i}(x)$ yields that $(f_{i}(x))_{i\in I}\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|f_{i}(x)\|_{(\mu,\theta,r^{\prime})}\leq\kappa$. This shows (ii). ∎
@@
Then by the norm equivalence from Theorem 3.1, we find a $C_{5}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.13:
$\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{5}\|(1-\mu_{i}g_{\lambda}(\mu_{i}))b_{i}\|_{(\mu,\theta,r)}$ (132)
$\leq C_{5}\sup_{i\in I}\left(\big|1-\mu_{i}g_{\lambda}(\mu_{i})\big|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (133)
@@
(ii). By the norm equivalence in Theorem 3.1, we find a $C_{6}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.13:
$\|[f_{P,\lambda,g}]_{\sim}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{6}\|\mu_{i}g_{\lambda}(\mu_{i})b_{i}\|_{(\mu,\theta,r)}$ (134)
$\leq C_{6}\sup_{i\in I}\left(|\mu_{i}g_{\lambda}(\mu_{i})|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (135)